how many bricks each measuring 25cm*11.25cm*6cm will be needed to build a wall 8m long,6m high, 22.5 m thick?
Answer by mananth(16949)

volume of wall / volume of each brick = number of bricks
Volume of wall = 8m*6m *22.5 m =1080 m^3
Volume of each brick = 25cm*11.25cm*6cm
convert to meters
0.25*0.1125*0.06=0.0016875 m^3
1 080 / 0.0016875 = 640,000 bricks